# The multimessenger view of Pulsar Timing Array black holes with the Horizon-AGN simulation

### v1

@@ 8 Conclusions
This paper is based on the results of the hydrodynamical cosmological simulation Horizon-AGN (Dubois et al., 2014a), which was used to extract MBHBs and investigate their properties. The hardening of MBH pairs identified in Horizon-AGN was followed through their interaction with the stellar and gaseous environment in post-processing, producing a catalog of inspiralling MBHBs emitting in the PTA band. Each MBHB in the simulation was assigned an initial orbital eccentricity motivated by the numerical results of Roedig et al. (2011). In addition, we have also considered a purely circular population of MBHBs. Based on these results, we have produced 2 0002\,000 universe realizations reflecting the stochasticity or cosmic variance, inherent to the binary population. The GW signal from an individual MBHB is almost monochromatic over the PTAs observation time for circular binaries. The power of the GW signal from eccentric MBHBs is distributed over harmonics of the orbital (mean anomaly) frequencies with a peak shifting to higher modes as eccentricity increases. The incoherent superposition of GWs from multiple MBHBs creates a stochastic signal - referred to as a GWB - at low frequencies. Particularly bright sources that stand above the GWB could be potentially resolved (CW candidates). We have considered two PTAs: one based on the recent EPTA DR2new dataset and one that reflects our projection to future SKA observations. In this manuscript, we study the properties of MBHBs forming the GWB and CW candidates. The main findings could be summarized as follows.
@@
  • Generation of the GW strain from the population of MBHBs. We have considered two approaches to building predictions of the GWB in the PTA band: GE and Population based. We claim that Population approach is more reliable and closer to the actual observations. The GE approach is one of the most used in the literature (see e.g. Rosado et al., 2015; Agazie et al., 2023c; EPTA Collaboration et al., 2024a). It converts the characteristic GW strain from the population of MBHBs into a timing residuals power spectral density using the pulsar response function averaged over sky position and polarization. However, this approach has several limitations. First, it does not account for either the finite duration of observations or the finite sensitivity of the PTA, both of which can significantly impact the inference of the GW background spectrum. Second, the GE approach underestimates the cosmic variance of the GW signal observed by PTAs, since it does not account for the reduction in the effective number of binaries contributing to the GW-induced timing residuals caused by the geometrical response of the pulsars. We show that the power spectra obtained using this method can deviate significantly from those inferred from PTA observations. The GW frequencies probed by PTAs are very close to the lowest frequency (defined as the inverse of the observation duration), and the effect of spectral leakage - although reduced by the timing model marginalization in the PTA data analysis - could potentially bias the GWB spectral inference towards a shallower power law with a higher amplitude at $f=1/\rm year$. This bias propagates in the astrophysical predictions. This should be carefully investigated for both circular and eccentric populations.
@@
  • Stochasticity of the combined GW signal at low frequencies. Based on 2 0002\,000 universe realizations we find that for both populations around an observing frequency of 3 nHz the GW signal is dominated by hundreds to thousands of binaries. Such high numbers confirm that, at low frequencies (111010 nHz), the GW signal is consistent with a Gaussian stochastic background. In addition to the GWB, we find that CW candidates with SNR >3>3 could be found in $4\%$ ($20\%$) of our universe realizations using the EPTA (SKA) sensitivity, assuming a 10 year-long observing duration. In most cases, their GW frequency is below 10 nHz.
@@
  • Characteristics of MBHBs at low GW frequencies. The largest contribution to the GWB at 3 nHz (low frequencies) comes from heavy binaries with the chirp mass in the range $10^{8.5}$$10^{9.5}\,\,{\rm M_{\odot}}$ and within the redshift layer $0.05<z<1.2$. They are hosted primarily in galaxies and halos only relatively more massive than the Milky Way, with a redshift distribution peaking around 0.30.30.40.4. The peak of the mass distribution for the eccentric population is slightly shifted toward higher-mass MBHBs. These sources are eccentric binaries at very low orbital frequencies that contribute to the 111010 nHz PTA band by emitting at high orbital harmonics. CW candidates are more massive and located at lower redshift. For the most part, they are hosted in group and cluster-like main halos with masses $10^{13}-10^{15}\;\,{\rm M_{\odot}}$ at low redshift, up to $z\sim 0.4$. A fraction of them are found to be hosted in lower mass subhalos.
@@
  • Electromagnetic properties of the MBHBs contributing to GW strain in the PTA band. We find that the binaries contributing most to the background, as well as the CW candidates, will typically appear as AGN rather than quasars, the latter corresponding to only 991616% of the CW candidates and just 33% of the background binaries. This is a consequence of their low Eddington ratio (below $10^{-2}$), which is typical of massive MBHs at low redshift. The host galaxies have low star formation rates, resulting from the time delay between the galaxy merger and the GW emission phase in the PTA band. However, we find that the CW candidates can outshine their host galaxies, particularly in the X-ray and radio bands, provided that their accretion rate satisfies $f_{\rm Edd}>10^{-3}$. Even for slow-accreting systems, under optimistic assumptions up to 8080% (3030%) of the CW candidates are brighter than their hosts in the radio (X-ray) band. This suggests that the radio band is the best suited for identifying multi-messenger MBHBs brighter than the host galaxies. Overall, since CW candidates are typically outshined by their hosts in the optical and near-infrared bands, searching for binaries using light curve variability in galactic nuclei is expected to be challenging. The existing MBHB candidates identified using such methods exhibit larger bolometric luminosities than the CW candidates extracted from our simulation. This indicates a potential selection bias in favor of more luminous, but rarer, MBHB candidates, or it suggests an alternative explanation for apparent periodicity. We have also noticed that CW candidates are mostly located in galaxy groups and clusters, suggesting that the CW sky location prior could follow the number density distribution of observed galaxies.
@@
  • The combined GW strain at high frequencies. At high frequencies, $f_{\rm GW}\ \sim 1/\rm year$ , the number of MBHBs contributing to GW strain drops to a few tens (several tens) for the circular (eccentric) population, indicating that the combined GW signal loses stochasticity. The GW strain from the eccentric population remains stochastic to somewhat higher frequency due to high harmonics contribution of MBHBs with low orbital frequency. This small number of binaries implies a larger cosmic variance, especially for the circular population, which is reflected in the properties of the contributing binaries at these frequencies. High-frequency MBHBs have lower chirp masses (in the range $10^{8}$$10^{9}\,\,{\rm M_{\odot}}$) compared to those contributing to the low-frequency GWB and are found at higher redshifts, typically in the range 0.40.411.
@@ Appendix A Relating merger density and number of inspiralling MBHBs
An intuitive way to obtain the characteristic strain, or equivalently the energy density of the GWB, is to add up the energy contribution from each inspiralling binary emitting in the PTA band by introducing ${\rm d}^{3}N_{\rm i}/{\rm d}z{\rm d}\@vec{\xi}{\rm d}\ln f_{\mathrm{orb}}$, the number of binaries in a layer of comoving volume encompassed between $z$ and $z+{\rm d}z$, with binary parameters in $\xi$, $\xi+{\rm d}\xi$] and orbital frequency in $\ln f_{\mathrm{orb}},\ln f_{\mathrm{orb}}+{\rm d}\ln f_{\mathrm{orb}}]$ (see Equation 27).
@@
The quantity $\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},\tau_{\rm c})$ is the rate of inspiralling binaries in redshift layer $z$ with parameters $\@vec{\xi}$ that reach a time to coalescence of $\tau_{\rm c}$. As a result, if one assumes that the population of inspiralling binaries is in a steady state on time scales comparable with the merger timescale of MBHBs emitting in the PTA band, at redshift $z$, this quantity does not depend on $\tau_{\rm c}$, such that $\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},\tau_{\rm c})=\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},0)$, which is exactly the rate of merging SMBHBs at redshift $z$, $\frac{{\rm d}^{2}N_{\rm m}}{{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}$. Hence, one can write,
@@
As a result, in a realistic universe realization, for a given merger $j$, a given observer would, on average, observe a number $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}})$ of such binary with orbital frequency in $\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2,\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2]$, that is
$\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}})$ $=\int{\rm d}z{\rm d}\@vec{\xi}\int_{\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2}^{\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2}{\rm d}\ln f_{\mathrm{orb}}\frac{{\rm d}^{2}n^{(j)}}{{\rm d}\@vec{\xi}{\rm d}z}\frac{{\rm d}z}{{\rm d}\tau_{\rm c}}\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}}\frac{{\rm d}V_{\rm c}}{{\rm d}z}$ (50)
$=\int{\rm d}z{\rm d}\@vec{\xi}\frac{{\rm d}^{2}n^{(j)}}{{\rm d}\@vec{\xi}{\rm d}z}\frac{{\rm d}z}{{\rm d}\tau_{\rm c}}\frac{{\rm d}V_{\rm c}}{{\rm d}z}\int_{\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2}^{\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2}{\rm d}\ln f_{\mathrm{orb}}\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}}$ (51)
@@
At the end, we recover Equation 29.
@@ Appendix B Optimization of the GWB computation
In order to have a realistic estimate of timing residuals induced by the GWB from a population of inspiralling MBHBs, we must sum up the contribution from each individual binary as presented in Section 5.1.2.
@@
Thus, we used some approximations in order to speed up the GWB timing residuals computation. First, we consider that for each of the catalog merger $j$, each of the $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}}^{(k)})$ binaries have the exact same orbital frequency $f_{\mathrm{orb}}^{(k)}$. Then, we consider two cases: (i) If the number of binaries drawn from the Poisson distribution with the mean $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}}^{(k)})$ is smaller than a threshold number $N_{\rm i}^{\rm(max)}=80$, we indeed compute each individual waveforms, assigning random orbital parameters and sky location to each source. The choice of $N_{\rm i}^{\rm(max)}$ results from a compromise between selecting a low value to reduce computational cost and a high value to ensure a valid approximation of the random walk described next.
(ii) If the number of binaries to model is greater than $N_{\rm i}^{\rm(max)}$, we use the following approximation. Let us introduce $N_{\rm i}^{(j)}(f_{\mathrm{orb}}^{(k)})$ ($N_{\rm i}^{\rm(max)}$ here) the number of binaries similar to the catalog binary $j$ and orbiting at $f_{\mathrm{orb}}^{(k)}$ that we have to model for a given universe realization. Using Equation 36, for each individual binary contribution $\tilde{r}_{\rm GW}^{(l)}(f)$, it is straightforward to see that their total contribution to the GWB timing residuals can be written as
@@
In order to avoid computing this potentially huge sum for all catalog mergers $j$ and orbital frequency bin $k$, it is useful to rewrite the sum as
@@
The key idea is to compute a fast estimate of the ratio of complex numbers appearing in the parentheses on the right-hand side. To achieve this, we approximate both sums as random walks in the complex plane, where each step has a random phase and an amplitude following the distribution of $\left|G_{\hat{k}}(\psi,\cos\iota)\right|$. Here, we assume a uniform distribution for the inclination angle, polarization angle, and binary sky location $\hat{k}$. The modulus of the ratio can thus be estimated by the ratio of the sums modulus standard deviations (their mean being zero). Following Mitra et al. (2021), it gives
@@
We can then assign a random complex phase to the modulus of Equation 56, to estimate the ratio in Equation 55. This way, the cost of the GWB timing residuals computation is greatly reduced without suffering much in accuracy, allowing us to compute thousands of universe realizations.
We note that instead of using this approximation, one can directly estimate Equation 54 in the case $N_{\rm i}^{(j)}\left(f_{\mathrm{orb}}^{(k)}\right)\geq N_{\rm i}^{\rm(max)}$ by invoking the central limit theorem. The total sum can be estimated by drawing two real numbers, $\hat{X}_{1}$ and $\hat{X}_{2}$, from two independent and identically distributed Gaussian variables $X_{1},X_{2}\sim\mathcal{N}\left(0,\sqrt{N_{\rm i}^{(j)}\left(f_{\mathrm{orb}}^{(k)}\right)\times\left\langle\left|G_{\hat{k}}\right|^{2}\right\rangle/2}\right)$, and computing $\hat{X}_{1}+i\hat{X}_{2}$. For the isotropic/uniform prior that we use here for $\hat{k},\cos\iota$ and $\psi$, we find $\langle\left|G_{\hat{k}}\right|^{2}\rangle=1/30$. We verified that the two methods yield highly consistent results, with deviations at the percentage level.

### v2

@@ 8 Conclusions
This paper is based on the results of the hydrodynamical cosmological simulation Horizon-AGN (Dubois et al., 2014a), which was used to extract MBHBs and investigate their properties. The hardening of MBH pairs identified in Horizon-AGN was followed through their interaction with the stellar and gaseous environment in post-processing, producing a catalog of inspiralling MBHBs emitting in the PTA band. Each MBHB in the simulation was assigned an initial orbital eccentricity motivated by the numerical results of Roedig et al. (2011). In addition, we also considered a purely circular population of MBHBs. Based on these results, we produced 2 0002\,000 Universe realizations reflecting the stochasticity or cosmic variance, inherent to the binary population. The GW signal from an individual MBHB is almost monochromatic over the PTAs observation times for circular binaries. The power of the GW signal from eccentric MBHBs is distributed over harmonics of the orbital (mean anomaly) frequencies with a peak shifting to higher modes as eccentricity increases. The incoherent superposition of GWs from multiple MBHBs creates a stochastic signal (referred to as a GWB) at low frequencies. Particularly bright sources that stand above the GWB could be potentially resolved (CW candidates). We considered two PTAs: one based on the recent EPTA DR2new dataset and one that reflects our projection of future SKA observations. In this paper, we study the properties of MBHBs forming the GWB and CW candidates. Our main findings can be summarized as follows.
@@
  • Generation of the GW strain from the population of MBHBs. We considered two approaches to building predictions of the GWB in the PTA band: GE-based and population-based. We claim that the population approach is more reliable and closer to the actual observations. The GE approach is one of the most commonly used in the literature (see e.g., Rosado et al., 2015; Agazie et al., 2023c; EPTA Collaboration et al., 2024a). It converts the characteristic GW strain from the population of MBHBs into a timing residuals power spectral density using the pulsar response function averaged over sky position and polarization. However, this approach has several limitations. First, it does not account for either the finite duration of observations or the finite sensitivity of the PTA, both of which can significantly impact the inference of the GW background spectrum. Second, the GE approach underestimates the cosmic variance of the GW signal observed by PTAs, since it does not account for the reduction in the effective number of binaries contributing to the GW-induced timing residuals caused by the geometrical response of the pulsars. We show that the power spectra obtained using this method can deviate significantly from those inferred from PTA observations. The GW frequencies probed by PTAs are very close to the lowest frequency (defined as the inverse of the observation duration), and the effect of spectral leakage (albeit reduced by the timing model marginalization in the PTA data analysis) could potentially bias the GWB spectral inference toward a shallower power law with a higher amplitude at $f=1/\rm year$. This bias propagates in the astrophysical predictions. This should be carefully investigated for both circular and eccentric populations.
@@
  • Stochasticity of the combined GW signal at low frequencies. Based on 2 0002\,000 Universe realizations we find that for both populations around an observing frequency of 3 nHz the GW signal is dominated by hundreds to thousands of binaries. Such high numbers confirm that at low frequencies (111010 nHz), the GW signal is consistent with a Gaussian stochastic background. In addition to the GWB, we find that CW candidates with S/N >3>3 could be found in $4\%$ ($20\%$) of our Universe realizations using the EPTA (SKA) sensitivity, assuming a ten year-long observing duration. In most cases, their GW frequency is below 10 nHz.
@@
  • Characteristics of MBHBs at low GW frequencies. The largest contribution to the GWB at 3 nHz (low frequencies) comes from heavy binaries with the chirp mass in the range $10^{8.5}$$10^{9.5}\,\,{\rm M_{\odot}}$ and within the redshift layer $0.05<z<1.2$. They are hosted primarily in galaxies and halos that are more massive than the Milky Way, with a redshift distribution peaking around 0.30.30.40.4. The peak of the mass distribution for the eccentric population is slightly shifted toward higher-mass MBHBs. These sources are eccentric binaries at very low orbital frequencies that contribute to the 111010 nHz PTA band by emitting at high orbital harmonics. CW candidates are more massive and located at lower redshift. For the most part, they are hosted in group and cluster-like main halos with masses $10^{13}-10^{15}\;\,{\rm M_{\odot}}$ at low redshift, up to $z\sim 0.4$. A fraction of them are found to be hosted in lower mass subhalos.
@@
  • Electromagnetic properties of the MBHBs contributing to GW strain in the PTA band. We find that the binaries contributing most to the background, as well as the CW candidates, will typically appear as AGN rather than quasars, the latter corresponding to only 991616% of the CW candidates and just 33% of the background binaries. This is a consequence of their low Eddington ratio (below $10^{-2}$), which is typical of massive MBHs at low redshift. The host galaxies are expected to have low star formation rates, resulting from the time delay between the galaxy merger and the GW emission phase in the PTA band. However, we find that the CW candidates can outshine their host galaxies, particularly in the X-ray and radio bands, provided that their accretion rate satisfies $f_{\rm Edd}>10^{-3}$. Even for slow-accreting systems, under optimistic assumptions up to 8080% (3030%) of the CW candidates are brighter than their hosts in the radio (X-ray) band. This suggests that the radio band is the best suited for identifying multi-messenger MBHBs brighter than the host galaxies. Overall, since CW candidates are typically outshined by their hosts in the optical and NIR bands, searching for binaries using light curve variability in galactic nuclei is expected to be challenging. The existing MBHB candidates identified using such methods exhibit larger bolometric luminosities than the CW candidates extracted from our simulation. This indicates a potential selection bias in favor of more luminous (but rarer) MBHB candidates; alternatively, it suggests that there must be another explanation behind the apparent periodicity. We also noticed that CW candidates are mostly located in galaxy groups and clusters, suggesting that the CW sky location prior could follow the number density distribution of observed galaxies.
@@
  • The combined GW strain at high frequencies. At high frequencies, $f_{\rm GW}\ \sim 1/\rm year$, the number of MBHBs contributing to GW strain drops to a few tens (several tens) for the circular (eccentric) population, indicating that the combined GW signal loses stochasticity. The GW strain from the eccentric population remains stochastic to a somewhat higher frequency due to the high harmonics contribution of MBHBs with low orbital frequency. This small number of binaries implies a larger cosmic variance, especially for the circular population, which is reflected in the properties of the contributing binaries at these frequencies. High-frequency MBHBs have lower chirp masses (in the range $10^{8}$$10^{9}\,\,{\rm M_{\odot}}$), compared to those contributing to the low-frequency GWB and are found at higher redshifts, typically in the 0.40.411 range.
Data availability
The data that support the findings of this study are openly available at https://doi.org/10.5281/zenodo.18391289.
@@ Appendix A Relating merger density and number of inspiralling MBHBs
An intuitive way to obtain the characteristic strain, or equivalently the energy density of the GWB, is to add up the energy contribution from each inspiralling binary emitting in the PTA band by introducing ${\rm d}^{3}N_{\rm i}/{\rm d}z{\rm d}\@vec{\xi}{\rm d}\ln f_{\mathrm{orb}}$, the number of binaries in a layer of comoving volume encompassed between $z$ and $z+{\rm d}z$, with binary parameters in $\xi$, $\xi+{\rm d}\xi$] and orbital frequency in $\ln f_{\mathrm{orb}},\ln f_{\mathrm{orb}}+{\rm d}\ln f_{\mathrm{orb}}]$ (see Eq. 27).
@@
The quantity $\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},\tau_{\rm c})$ is the rate of inspiralling binaries in redshift layer $z$ with parameters $\@vec{\xi}$ that reach a time to coalescence of $\tau_{\rm c}$. As a result, if one assumes that the population of inspiralling binaries is in a steady state on time scales comparable with the merger timescale of MBHBs emitting in the PTA band, at redshift $z$, this quantity does not depend on $\tau_{\rm c}$, such that $\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},\tau_{\rm c})=\frac{{\rm d}^{3}N_{\rm i}}{{\rm d}z{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}(z,\@vec{\xi},0)$, which is exactly the rate of merging SMBHBs at redshift $z$, $\frac{{\rm d}^{2}N_{\rm m}}{{\rm d}\@vec{\xi}{\rm d}\tau_{\rm c}}$. Hence, one can write
@@
As a result, in a realistic Universe realization, for a given merger $j$, a given observer would, on average, observe a number $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}})$ of such a binary with an orbital frequency in the range $\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2,\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2]$, namely,
$\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}})$ $=\int{\rm d}z{\rm d}\@vec{\xi}\int_{\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2}^{\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2}{\rm d}\ln f_{\mathrm{orb}}\frac{{\rm d}^{2}n^{(j)}}{{\rm d}\@vec{\xi}{\rm d}z}\frac{{\rm d}z}{{\rm d}\tau_{\rm c}}\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}}\frac{{\rm d}V_{\rm c}}{{\rm d}z},$ (50)
$=\int{\rm d}z{\rm d}\@vec{\xi}\frac{{\rm d}^{2}n^{(j)}}{{\rm d}\@vec{\xi}{\rm d}z}\frac{{\rm d}z}{{\rm d}\tau_{\rm c}}\frac{{\rm d}V_{\rm c}}{{\rm d}z}\int_{\ln f_{\mathrm{orb}}-\Delta\ln f_{\mathrm{orb}}/2}^{\ln f_{\mathrm{orb}}+\Delta\ln f_{\mathrm{orb}}/2}{\rm d}\ln f_{\mathrm{orb}}\frac{{\rm d}\tau_{\rm c}}{{\rm d}\ln f_{\mathrm{orb}}},$ (51)
@@
Finally, at the end of this calculation, we recover Eq. 29.
@@ Appendix B Optimization of the GWB computation
To have a realistic estimate of timing residuals induced by the GWB from a population of inspiralling MBHBs, we must sum up the contribution from each individual binary as presented in Sect. 5.1.2.
@@
Thus, we used some approximations in order to speed up the GWB timing residuals computation. First, we consider that for each of the catalog merger $j$, each of the $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}}^{(k)})$ binaries has the exact same orbital frequency $f_{\mathrm{orb}}^{(k)}$. Then, we consider two cases: (i) if the number of binaries drawn from the Poisson distribution with the mean $\langle N_{\rm i}^{(j)}\rangle(f_{\mathrm{orb}}^{(k)})$ is smaller than a threshold number $N_{\rm i}^{\rm(max)}=80$, we indeed compute each individual waveform, assigning random orbital parameters and sky location to each source. The choice of $N_{\rm i}^{\rm(max)}$ results from a compromise between selecting a low value to reduce computational cost and a high value to ensure a valid approximation of the random walk described next.
(ii) If the number of binaries to model is greater than $N_{\rm i}^{\rm(max)}$, we can use the following approximation. We start by introducing $N_{\rm i}^{(j)}(f_{\mathrm{orb}}^{(k)})$ ($N_{\rm i}^{\rm(max)}$ here), which is the number of binaries similar to the catalog binary, $j$, and orbiting at $f_{\mathrm{orb}}^{(k)}$ that we have to model for a given Universe realization. Using Eq. 36, for each individual binary contribution $\tilde{r}_{\rm GW}^{(l)}(f)$, it is straightforward to see that their total contribution to the GWB timing residuals can be written as
@@
To avoid computing this potentially huge sum for all catalog mergers $j$ and orbital frequency bin $k$, it is useful to rewrite the sum as
@@
The key idea is to compute a fast estimate of the ratio of complex numbers appearing in the parentheses on the right-hand side. To achieve this, we approximate both sums as random walks in the complex plane, where each step has a random phase and an amplitude following the distribution of $\left|G_{\hat{k}}(\psi,\cos\iota)\right|$. Here, we assume a uniform distribution for the inclination angle, polarization angle, and binary sky location $\hat{k}$. The modulus of the ratio can thus be estimated by the ratio of the sums of the modulus standard deviations (their mean being zero). Following Mitra et al. (2021), it gives
@@
We can then assign a random complex phase to the modulus of Eq. 56, to estimate the ratio in Eq. 55. This way, the cost of the GWB timing residuals computation is greatly reduced without suffering much in accuracy, allowing us to compute thousands of Universe realizations.
We note that instead of using this approximation, one can directly estimate Eq. 54 in the case $N_{\rm i}^{(j)}\left(f_{\mathrm{orb}}^{(k)}\right)\geq N_{\rm i}^{\rm(max)}$ by invoking the central limit theorem. The total sum can be estimated by drawing two real numbers, $\hat{X}_{1}$ and $\hat{X}_{2}$, from two independent and identically distributed Gaussian variables $X_{1},X_{2}\sim\mathcal{N}\left(0,\sqrt{N_{\rm i}^{(j)}\left(f_{\mathrm{orb}}^{(k)}\right)\times\left\langle\left|G_{\hat{k}}\right|^{2}\right\rangle/2}\right)$, and computing $\hat{X}_{1}+i\hat{X}_{2}$. For the isotropic/uniform prior that we use here for $\hat{k},\cos\iota$ and $\psi$, we find $\langle\left|G_{\hat{k}}\right|^{2}\rangle=1/30$. We verified that the two methods yield highly consistent results, with deviations at the percentage level.